OpenStudy (anonymous):

A particle moves along a straight line with equation of motion s = f(t), where s is measured in meters and t in seconds. Find the velocity and speed when t = 8. f(t) = t^-1 - t Its my last attempt and i cant get it :(

OpenStudy (anonymous):

do u think it will be -65/64 and 65/64

OpenStudy (anonymous):

i'm chating with the guy about your problem now

OpenStudy (bahrom7893):

f(t) = t^-1 - t V(t) = -1*t^(-2) - 1 V(8) = (-1/64) - 1 = -65/64 Speed at t = 8 => |-65/64| = 65/64, so yes, ur right!